OpenStudy (anonymous):

A machine gun fires 25 g bullets at the rate of 600 bullets per minute with a speed of 200 m/s. Calculate the force required to keep the gun in position.

OpenStudy (anonymous):

600 bullets per minute = 600/60 = 10 bullets per second F = n * m * v = 10 * 200 * 25 * 10^-3 = 2N

OpenStudy (anonymous):

But force is rate of change of momentum right?

OpenStudy (anonymous):

sorry F=50N

OpenStudy (anonymous):

He gave you in fact Change of momentum/ per 1 sec

OpenStudy (anonymous):

My calcu was wrong..))

OpenStudy (anonymous):

I didn't get you @Mikael.

OpenStudy (anonymous):

See he calculated for U how much Momentum is given to ALL the bullets flying-out in the period of 1 second . IT IS the change in momentum of the gun too - with a minus sign of course

OpenStudy (anonymous):